OpenStudy (anonymous):

I have to write the equation of each line Through (-4, 3) and with slope =0 is it y-y=m (x-x)?

hero (hero):

Since m = 0, x = -4 and y = 3, replace those values in the formula like so: y-y=m (x-x) y - 3 = 0(x+4) Anything times zero equal zero: y - 3 = 0 y = 3

OpenStudy (anonymous):

ty

hero (hero):

See, once you actually post the full sentence, I am able to help you :P